OpenStudy (phebe):

14. Use the Zero Product Property to solve the equation: -3m(2m – 5) = 0

jimthompson5910 (jim_thompson5910):

The idea of the Zero Product Property is that if A*B = 0 then A = 0 or B = 0

jimthompson5910 (jim_thompson5910):

So how would you use that idea in this case?

OpenStudy (phebe):

yess ik that thanx thoee i juss need help with the question

OpenStudy (phebe):

do imultiply

jimthompson5910 (jim_thompson5910):

think of A as -3m think of B as 2m-5

OpenStudy (acxbox22):

when multiplied, the product is zero. this means that one factor is zero. so you just set both factors equal to zero and solve -3m=0 and 2m-5=0

OpenStudy (phebe):

okk

OpenStudy (phebe):

so do u know the answer

OpenStudy (acxbox22):

just solve the 2 equations

OpenStudy (phebe):

ok hold on k

OpenStudy (phebe):

i still don't know if my answer is ruight

OpenStudy (phebe):

can u tell me the answer

jimthompson5910 (jim_thompson5910):

tell us what you get so we can check

OpenStudy (phebe):

ok hold on

OpenStudy (phebe):

i got 3

OpenStudy (phebe):

@yahman

OpenStudy (anonymous):

wat u need

OpenStudy (acxbox22):

-3m=0 and 2m-5=0 -3m=0 divide both sides by -3 (hint: 0 divided by anthing is 0) 2m-5=0 2m=5 divide both sides by 5

OpenStudy (phebe):

ok i got that @acxbox22 is that all i need to do

OpenStudy (phebe):

@yahman do u know ho wwot do this

OpenStudy (anonymous):

ya wait hmmmmmmmmmmm

OpenStudy (acxbox22):

yeah, plus you are going to have 2 solutions , 1 from each equation

OpenStudy (phebe):

ok @yahman nd @acxbox22

OpenStudy (anonymous):

Get Zero on one side of the equal sign Factor the equation Split the equation up Solve each equation

OpenStudy (phebe):

ok

OpenStudy (anonymous):

four steps

OpenStudy (phebe):

yeass ik

OpenStudy (anonymous):

Basic Examples

OpenStudy (phebe):

yea thas ok forget this question i'll jus guess on it k

OpenStudy (phebe):

thanc anywayss

OpenStudy (anonymous):

Solve (x + 3)(x - 5) = 0 -Zero is on one side of the equal sign -The equation is already factored -Split each set of parentheses into its own equation x + 3 = 0 or x - 5 = 0 x = -3 or x = 5 Solve each equation {-3, 5}

OpenStudy (anonymous):

this is the example

OpenStudy (phebe):

ok

OpenStudy (anonymous):

see my pic pheb

OpenStudy (phebe):

yess

OpenStudy (anonymous):

post the qs

OpenStudy (phebe):

ok ima put itin another post then tag u k

OpenStudy (anonymous):

kk